OpenStudy (anonymous):

302degrees f is how many Celsius

OpenStudy (anonymous):

150

OpenStudy (amistre64):

if we derive a formula from 2 points so that we have a linear relationship between f and c; itd be a neato

OpenStudy (anonymous):

what is a neato?? And you wanted to find a correlation between both of these units ??

OpenStudy (amistre64):

f,c 212,100 -(32 , 0) -------- 180,100 ; slope = 10/18 = 5/9 \[C(f)=\frac{5}{9}f-\frac{5(32)}{9}+0\] \[C(f)=\frac{5}{9}f-\frac{160}{9}\] \[C(f)=\frac{5f-160}{9}\] maybe?

OpenStudy (amistre64):

C(320) = [5(320)-160]/9 = (1600-160)/9 = 1440/9 = 160

OpenStudy (amistre64):

helps if i can read "302" and not 320 :) but the steps are the same, just plug in the right number :)

OpenStudy (amistre64):

.. 150